OpenStudy (anonymous):

Find the slope of the line containing the points (-2, 7) and (1, 2) Answer -1/3 -5/3 -3/5 -1/2

OpenStudy (anonymous):

use formula slope =(y2-y1)/(x2-x1)

OpenStudy (anonymous):

So option B

OpenStudy (anonymous):

(2-7)(1--2) = (2-7)(1+2) = -5/3